What annual rate of return is earned on a $5,000 investment when it grows to $9,000 in five years? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Here we will use the following formula:

FV = PV (1+R%)n

where, FV= Future Value , PV = Present value, R% is required rate of return and n = time period.

Now, putting the values into the above equation, we get,

9000 = 5000 (1+ R%)5

9000 / 5000 = (1+R%)5

1.8 = (1+R%)5

(1.8)1/5 = 1+R%

1.12474611314 = 1 +R%

R% = 1.12474611314 -1

R% = 0.12474611314

R = 12.47%